Suppose that a typical firm in a monopolistically competitive industry faces a demand curve given by:


q = 60 − (1/2)p, where q is quantity sold per week.


The firm’s marginal cost curve is given by: MC = 60.

  1. How much will the firm produce in the short run?
  2. What price will it charge?

In addition to providing the quantitative answers for the question, please also describe the approach you used to arrive at your conclusions.


1
Expert's answer
2021-12-15T22:35:00-0500

(a)

"q=60-\frac{1}{2}p"

"\frac{1}{2}p=60-q"

"p=120-2q"

"TR=(120-2q)q"

"TR=120q-2q^2"

"MR=120-4q"

"MC=60"

In the short run, the firm should produce quantity for which profit is maximized,

We thus take:

"MC=MR"

"60=120-4q"

"-60=-4q"

"q=15"

(b)

"P=120-2q"

"=120-2(15)=90."